OpenStudy (anonymous):

what is the greatest value if z=1/2 z^-6z^4

OpenStudy (anonymous):

put the value it becomes it greatest as whole expression becomes constant

OpenStudy (anonymous):

please explain

OpenStudy (anonymous):

z^-6+4=z^-2 putting z=1/2 we have 1/2^-2 =a constant value so, its only value =greatest=least